OpenStudy (anonymous):

The equation of the quadratic function with roots 6 and 10 and a vertex at (8, 2) is f of x equals 1/2x^2 plus 8x-30. True or false? Medals given(:

OpenStudy (anonymous):

8=D

OpenStudy (anonymous):

give me my medal lol

OpenStudy (anonymous):

Um..

OpenStudy (campbell_st):

well lots of ways to look at it... you know the roots so f(x) = a(x -6)(x -10) which is \[f(x) = a(x^2 - 16x + 60)\] looking at the given value of a, its a = 1/2 so substiute the vertex x = 8 and y = 2 so see if a = 1/2 so 2 = 1/2(8^2 -16(8) + 60) 2 = 1/2(-4) so does 2 = -2 thats what you need to decide

OpenStudy (anonymous):

kiiid freaking man, i love u, anyways @beccaboo333 can help with this problem

OpenStudy (anonymous):

So it's false?

OpenStudy (anonymous):

@beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333 @beccaboo333

OpenStudy (campbell_st):

that would be my answer

OpenStudy (anonymous):

Next question Find the equation of the quadratic function with zeros 10 and 14 and a vertex at (12, -8).

OpenStudy (campbell_st):

well f(x) =a(x -10)(x -14) substitute (12, -8) to find a

OpenStudy (anonymous):

y = -2x2 − 48x + 280